OpenStudy (anonymous):

Solve for n and select the correct answer. -2(7n + 10) = 7(4 – 2n) n = 28 n= 48 No solution All real solutions

OpenStudy (anonymous):

-2(7n + 10) = 7(4 – 2n) Distributing -2 and 7, -14n - 20 = 28 - 14n 0n = 48 no solution